Find v, and i, in the circuit shown below: 102 3A A 50 V + 212 + + 4i. 4,
Answer #1

10JL -> Vx 52 50V 4 lx So l:- APply kul for +the SuperMesh, 50- t0i, -53 - 4°2=0 ( x=) 50=14i, +5°2 Apply kel at nade A, i+3+Substitute 3 in 2, i+3+(2(-i2)) = i2 1.5° +3=1.5°2 -1.Sit1:5°2=3->@ (4) solving O &@, we get := 2-105 AMP, 2=4.l05 Amp From ®